OpenStudy (anonymous):

hi can someone help me please because i need help well this is the problem: the sum of the digits of a two digit number is 12 times the tens digit. find the number.

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

:,( i need help cuz i dont get it

OpenStudy (anonymous):

supervisor plz help

OpenStudy (anonymous):

12 times the tens digit... So the sum of the digits has to be at least 12... I do not think this is the right question... Do you mean three digits?

OpenStudy (anonymous):

no just two digits thats what it says on my paper

OpenStudy (anonymous):

Reason being: Even if the tens digit is 1, then the sum of the digits would have to be 12 - in other words, the ones digit would have to be 11... Impossible?

OpenStudy (anonymous):

00, 01, 02, 03, 04, 05, 06, 07, 08, 09 all work... That's the witty response, though.

OpenStudy (anonymous):

!! Wrong, I didn't see it had to be equal to 12 times the value... I don't think there is an answer.

OpenStudy (anonymous):

really ok how about this one cuz i got stuck on it two: the sum of the digit of a two-digit number is 12.if 15 is added to the number,the result is 6 times the units digits. find the number

OpenStudy (anonymous):

That seems plausible: Let x = tens digit; y = units digit Thus, 10x + y = the number x + y = 12 10x + y + 15 = 6y y = 2x + 3 (simplifying) Plug that into x + y = 12 x + 2x + 3 = 12 3 x = 9 x = 3 y = 9 39